# Learning on LLM Output Signatures for Gray-Box Behavior Analysis

### v2

@@ 3.2 Our Approach
4 Generalization of Previous Approaches
Here, we demonstrate that LOS-Net generalizes several leading existing methods through specific weight configurations. This ensures that our architecture can theoretically match these methods –while significantly outperforming them in practice, as shown in our experiments. Proofs are enclosed in Appendix A. As already mentioned, prior research has introduced various methods for analyzing LLMs based on their output probabilities [14, 22, 49, 18], with many approaches focusing on the ATPs. We note that many of these methods assume the form of statistics calculated over the whole sequence processed by the LLM. Recent, more sophisticated approaches aggregate these probabilities only for some of the tokens in the sequence, dynamically chosen based on features computed on the set of ATPs [41, 58], as we illustrate below.
Motivating example: Min-K%  Shi et al. [41]. Min-Kmakes predictions on an input text $\vec{s}$ based on a score $R$ calculated as the average of the smallest $K$log-probs: $R(\vec{s})=\frac{1}{|M|}\sum_{i\in M}\log(p_{i})$, with $M=\{i\mid p_{i}<\text{perc}(\mathbf{p},K)\}$ being the set of token indices whose probabilities are in the first $K$-th percentile of $\mathbf{p}$. We note that it is instructive to rewrite the scoring equation as:
$R(\vec{s})=\sum_{i=1}^{|\vec{s}|}\overbrace{\frac{\log(p_{i})}{\left\lceil\frac{K}{100}\cdot|\vec{s}|\right\rceil}}^{\text{token-wise score}}\cdot\underbrace{\mathbb{I}\big{(}\overbrace{p_{i}}^{\text{confidence}}<\overbrace{\text{perc}(\mathbf{p},K)}^{\text{adaptive threshold}}\big{)}}_{\text{gating}}.$ (4)
@@ 4 Generalization of Previous Approaches
$F(\mathbf{X},\mathbf{p})_{i}$ =$=\begin{cases}g(\mathbf{X}^{\prime},\mathbf{p})_{i},&\text{if }\kappa(\mathbf{X}^{\prime},\mathbf{p})_{i}\geq T(\mathbf{X}^{\prime},\mathbf{p}),\\ 0,&\text{otherwise},\end{cases}\qquad R(\mathbf{X},\mathbf{p})$ $=\sum_{i=1}^{N}F(\mathbf{X},\mathbf{p})_{i},$ (5)
where $\mathbf{X}^{\prime}$ is the sorted version of $\mathbf{X}$, as per Equation 1. The family of GSF is flexible enough to capture previously proposed gray-box methods, as we show in the following:
Proposition 4.1 (GSFs capture known baselines).
Let $\mathcal{B}$ be the set of scoring functions implemented by the Min/Max/Mean aggregated probability methods [14, 22, 49, 18] for HD, as well as the MinK[41] and MinK%++ [58] methods for DCD. For any scoring function $f\in\mathcal{B}$, there exists a choice of functions $\kappa,T,g$ such that the GSF $R$ in Equation 5, implements $f$.
It is easy to see, e.g., how MinKis implemented as a GSF. For a sequence length of $N$, it suffices to choose: $T(\mathbf{X}^{\prime},\mathbf{p})=-\text{perc}(\mathbf{p},K)=-\big{(}\text{sort}(\mathbf{p})_{\left\lceil\frac{K}{100}\cdot N\right\rceil}\big{)},\quad\kappa(\mathbf{X}^{\prime},\mathbf{p})=-\mathbf{p},\quad g(\mathbf{X}^{\prime},\mathbf{p})=\frac{\log\mathbf{p}}{\left\lceil\frac{K}{100}\cdot N\right\rceil}.$ Refer to Appendix A for more details on how other baselines are implemented.
LOS-Net can approximate GSFs and implement known baselines. As the following results show, our LOS-Net architecture is theoretically justified from an expressiveness standpoint. We start by showing that it can, in fact, approximate virtually all GSFs of interest.
Proposition 4.2 (LOS-Net can approximate Equation 5).
@@
To prove this result, we build on existing universality results on approximating continuous functions with Transformers [57], showing that our (generally non-continuous) target functions can be approximated by continuous functions. Importantly, Proposition 4.2 implies that, as long as the LOS space of interest lies within a compact domain [^3] This is inherently satisfied when using probabilities; or via clamping in the case of logits or log-probs., our model can approximate the general GSF in Equation 5 of LOSs for any LLM under mild conditions on $\kappa$, $T$, and $g$, potentially generalizing across LLMs (as our result considers a predefined LOS domain). In Section 5 we show that our trained models can indeed be applied successfully out-of-the-box on LOSs from different LLMs. Note that Proposition 4.2 cannot be generally extended to $L_{\infty}$ due to the discontinuity of GSFs.
The practical relevance of Proposition 4.2, is underscored by the following:
Corollary 4.3 (Approximation of Baselines by LOS-Net).
Our architecture, as defined in Equation 3, can arbitrarily well approximate, in the $L_{1}$ sense, any of the baseline methods in $\mathcal{B}$ when operating on context and token-vocabulary of, resp., maximal sizes $N_{\text{max}}$ and $V_{\text{max}}$.
The above states that well-established, successful baselines from the literature (see class $\mathcal{B}$ in Proposition 4.1) can be approximated by LOS-Net. The proof follows from Propositions 4.1 and 4.2.
5 Experiments
We assess various aspects of learning with LOS via the following questions: (Q1) Is learning on LOS an effective approach for addressing key tasks such as DCD and HD? Does it outperform baselines? (Sections 5.2 and 5.1); (Q2) Does our model exhibit transfer capabilities across LLMs and datasets, suggesting the emergence of universal patterns in LLM behavior from the perspective of the LOS? (Section 5.3); (Q3) How important is $\mathbf{X}$ in the pair $(\mathbf{X},\mathbf{p})$, as it is often overlooked? And how impactful is the choice of the slicing parameter $K$ in Equation 1? (Section B.6). In the following, we present our main results, and refer to Appendix B for additional experimental results and details.
General setup.
Our experiments focus on the two tasks of DCD and HD, with hyperparameter $K$ fixed at $1000$ unless stated otherwise (see Equation 1). In Section B.6, we show that our model is robust to variations in $K$. To align with prior work, we use datasets and LLMs from [41, 58] for DCD and [35] for HD, where we also experiment with an additional LLM (Qwen-2.5-7b-Instruct [53]). Further details are in subsequent sections. We use the area under the ROC curve (AUC) to evaluate HD and DCD, a standard metric in this domain [35, 41, 58], which measures the balance between sensitivity and specificity. We conduct each experiment across three different random seeds (when applicable) and report the mean along with the standard deviation of the results. All LOS-Net experiments were conducted using the PyTorch [36] framework on a single NVIDIA L-40 GPU.
Newly introduced learning-based baselines. In addition to task-specific baselines, we also introduce two novel learning-based baselines to appreciate the contribution of the TDS: ATP+R-MLP, ATP+R-Transf.. Specifically, we ablate information about the TDS and only process the ATP and rank information with an MLP or Transformer backbone. Formal definitions are in Section B.4.
5.1 Hallucination Detection
We follow the setup of Orgad et al. [35]. The objective is to predict whether an LLM-generated response to a given input prompt is correct or not. We frame the task within a gray-box setting, i.e., we assume no access to the LLM’s internals. We assume no access to external resources – such as auxiliary LLMs (as in [47]), mulitple queries/generations (as in [35, 23]), or additional contextual cues like token-level annotations ([35]). These approaches exhibit a significantly larger detection latency compared to LOS-Net, problematic in online applications (refer to Section 5.4 for more detailed run-time analyses and comparisons).
Datasets and LLMs. Following Orgad et al. [35], we use three datasets spanning various domains and tasks: HotpotQA without context [54], IMDB sentiment analysis [30], roles in Movies [35]. Details regarding the annotation process, splits and dataset sizes are in Section B.5.1. As the target LLMs, coherently with Orgad et al. [35], we use Mistral-7b-instruct-v0.2 [21] (Mis-7b) and LLaMa3-8b-instruct [46] (L-3-8b), and further experiment with Qwen-2.5-7b-Instruct [53] (Q-2.5-7b).
HD Baselines. (1) Aggregated probabilities/logits: Previous studies [14, 22, 49, 18] simply aggregate output token probabilities or logits to score LLM confidence for error detection. These aggregations operate mean/max/min pooling over the ATP. We refer to them as Logit/Probas-mean/min/max; (2) P(True): [22] found that LLMs show reasonable calibration in assessing their own correctness.
Table 1: Test AUCs for HD over Mis-7b and L3-8b (bold: best method, underlined: second best).
@@ 5.1 Hallucination Detection
Results. Table 1 presents a comprehensive summary of results on the LLMs considered in Orgad et al. [35] (Mis-7b and L-3-8b). These clearly demonstrate that LOS-Net outperforms all baselines across all six dataset/LLM combinations, often by a significant margin. For instance, on the IMDB dataset, LOS-Net achieves an AUC improvement of around 32 units over the best baseline for Mis-7b and 17 over the best baseline for L-3-8b. Our results further indicate that ATP learning-based baselines consistently underperform compared to LOS-Net, underscoring the critical role of the TDS, $\mathbf{X}$. Our ATP-based learnable baselines still outperform non-learnable methods in most cases, suggesting that a learning approach relying exclusively on ATP can still be a viable solution in certain scenarios. Results on Q-2.5-7b are consistent the above findings, and are deferred to Section B.7.
@@ 5.2 Data Contamination Detection
The goal in DCD is to determine if an LLM was trained on specific data. The raw dataset $D=\{q_{i},y_{i}\}_{i=1}^{\ell}$ contains ${\ell}$ text samples, where $q_{i}$ represents the text and $y_{i}$ indicates whether it was part of the training data. DCD is often framed as a Membership Inference Attack (MIA) [42, 32, 41].
Datasets and LLMs. We use three datasets to assess DCD, specifically: WikiMIA-32 and WikiMIA-64 [41], as well as BookMIA [41]. The WikiMIA-32 and -64 datasets contain excerpts from Wikipedia articles, consisting of, resp., 32 and 64 words. The distinction between contaminated and uncontaminated data is determined by timestamps. As in [41, 58], we attack Mamba-1.4b [13] (M-1.4b), LLaMa-13b/30b [46] (L-13b/30b), Pythia-6.9b [4] (P-6.9b). BookMIA is a dataset of book excerpts. Positive members correspond to books known to be well memorized by certain OpenAI models [11], or otherwise known to (partly) be in pretraining corpus of other open-source LLMs [1]. Non-members include excerpts from books released after 2023, necessarily absent from the pretraining corpus of the last ones. Interestingly, this dataset allows us to test LOS-Net’s DCD capability in a realistic scenario akin to copyright-infringement detection. We thus propose a new split that ensures all excerpts from the same book always appear either in the training or test split (and never in both). Details are enclosed in Section B.5.2. We attack LLMs considered in [1]: LLaMa-13b/30b [46] (L-13b/30b), Pythia-6.9b/12b[4] (P-6.9b/12b).
Table 2: Test AUCs on BookMIA. ‘P’: Pythia, ‘L’: LLaMa-1 (bold: best method, underlined: second best, pink: reference-based).
@@
DCD Baselines. The Loss approach [55] directly uses the loss value as the detection score. The Reference (Ref) method [9] calibrates the target LLM’s perplexity leveraging a similar reference model known or supposed not to have memorized text of interest [^4] For example for Pythia-12b, a valid reference LLM would be the smaller Pythia-70M.. Both Zlib and Lowercase [9] are also reference-based methods: they utilize zlib compression entropy and lowercased text perplexity as reference for normalization. Lastly, Min-K[41] and Min-K%++ [58] are reference-free methods, which examine token probabilities and average a subset of the minimum token scores, or a function thereof, over the input. For these baselines, we select their hyperparameters by maximizing performance on the validation set(s).
Results on BookMIA.
Refer to Table 2. LOS-Net attains exceptional results, largely surpassing other reference-free approaches. Among these last ones, ours is the only method that can match or outperform even the reference-LLM-based baselines. Importantly, crucial to such strong reference-free performance is to access, even partially, the TDS: our ATP-based learnable methods – which only process features for the actual sequence tokens – incur indeed significant performance drops.
Results on WikiMIA. Due to space limits, the full results on are provided in Section B.8, Table 5. LOS-Net consistently surpasses all baselines across all eight combinations of LLMs and datasets. The second-best method is MinK%++, followed by MinK%, consistent with the findings of [58].
@@ 5.3 Generalization to Other LLMs and Datasets
Figure 3: BookMIA zero-shot generalization (bold: outperforms ref-free baselines, : outperforms ref-based ones).
We further study the possibility to apply our models to settings different from those they were originally trained on. We focus on two variables: datasets and LLMs. Generalization across datasets was originally studied in [35] within the scope of HD and in the context of white-box setups; their relevance lies in the fact that non-trivial dataset generalization would potentially suggest a ‘universal truthfulness’ representation encoded in the internal states and/or outputs of an LLM [35, 31, 43]. Inspecting transfer across LLMs is, to the best of our knowledge, still unexplored. This study is important for learning-based approaches in applications such as copyright-infringement detection, where ground-truth labels may be scarce.
@@
Zero Shot Cross-LLM Generalization Capabilities in DCD. We assess our model’s ability to detect DC in target LLMs that were unseen during training. Using the BookMIA benchmark and the setup described in Section 5.2, we evaluate our model directly across different LLMs without any fine-tuning. This setup is relevant in cases where contamination information is not yet available for newly released LLMs. The results are presented in the heatmap shown in Figure 3. We observe strong transferability: in 10/12 cases, our model achieves the best performance among reference-free approaches, highlighted in bold in Figure 3. Interestingly, in 3/12 cases, LOS-Net (which is reference-free) even surpasses reference-based baselines, as indicated via a superscript of . We also observe particularly strong transfer across differently sized LLM architectures within the same family and highlight the surprising positive transfer from the largest LLaMa to Pythia models.
@@
Transfer Learning across LLMs and Datasets for HD. Differently than DCD, where zero-shot application of LOS-Net was successful, for HD we observed non-trivial generalization in the zero-shot setup, however, not sufficient to surpass the simple probability-based techniques. This led us to investigate LOS-Net capabilities in a transfer learning setting, in which we conduct a rapid fine-tuning procedure on all possible LLM/datasets combinations. Specifically, we perform a 10-epoch fine-tuning on the target LLM/dataset (as opposed to 300+ epochs in our standard setting). This process was measured to take less than a minute. We benchmark the fine-tuned model against two baselines. First, to test for successful transfer, we compare with a LOS-Net trained from scratch under an identical setup (i.e., 10 epochs). Second, we contrast the fine-tuned model with the best-reported non-learnable baseline. The test AUC of our fine-tuned LOS-Net’s are in Figures 4 and 5. Superscript ‘’ indicates the fine-tuned LOS-Net is better than a counterpart trained from scratch in the same setting, bold indicates it outperforms the best non-learnable method.
@@
Discussion. First, LOS-Net exhibits solid transferability in both scenarios. The finetuned models consistently outperform their counterparts trained from scratch: 16/18 cases in both the cross-LLM (Figure 4) and cross-dataset setups (Figure 5) – see ‘’ on the off-diagonal entries. This highlights a generally positive transfer of LOS-Net’s learned representations across datasets and LLMs, and underscores the suitability of LOS as a data type in capturing generalizable patterns in LLM behaviors. Second, from a practical perspective, we find that LOS-Net outperforms the best baseline in 15/18 cases for both the cross-LLM (Figure 4) and cross-dataset (Figure 5) scenarios – see bold on the off-diagonal entries. Focusing on the IMDB dataset, when training on L-3-8b and testing on Mis-7b (Figure 4), our model substantially gains around 27 AUC units over the best baseline. This result underscores the possibility of transferring across LLMs. A similar trend is observed in the cross-dataset setup (Figure 5): on Mis-7b, when training on HotpotQA or Movies and testing on IMDB, our model achieves a notable improvement of around 30 AUC units compared to the best baseline.

### v3

@@ 4 Generalization of Previous Approaches
$F(\mathbf{X},\mathbf{p})_{i}$ $=\begin{cases}g(\mathbf{X}^{\prime},\mathbf{p})_{i},&\text{if }\kappa(\mathbf{X}^{\prime},\mathbf{p})_{i}\geq T(\mathbf{X}^{\prime},\mathbf{p}),\\ 0,&\text{otherwise},\end{cases},$
$R(\mathbf{X},\mathbf{p})$ $=\sum_{i=1}^{N}F(\mathbf{X},\mathbf{p})_{i},$ (4)
where $\mathbf{X}^{\prime}$ is the sorted version of $\mathbf{X}$, as per Equation˜1. The family of GSF is flexible enough to capture previously proposed gray-box methods, as we show in the following:
Proposition 1 (GSFs capture known baselines).
Let $\mathcal{B}$ be the set of scoring functions implemented by the Min/Max/Mean aggregated probability methods [16, 24, 49, 20] for HD, as well as Loss [55], the MinK% [43] and MinK%++ [58] methods for DCD. For any scoring function $f\in\mathcal{B}$, there exists a choice of functions $\kappa,T,g$ such that the GSF $R$ in Equation˜4, implements $f$.
It is easy to see, e.g., how MinK% is implemented as a GSF [^3] For a sequence length of $N$, it suffices to choose: $T(\mathbf{X}^{\prime},\mathbf{p})=-\text{perc}(\mathbf{p},K)=-\big(\text{sort}(\mathbf{p})_{\left\lceil\frac{K}{100}\cdot N\right\rceil}\big),\quad\kappa(\mathbf{X}^{\prime},\mathbf{p})=-\mathbf{p},\quad g(\mathbf{X}^{\prime},\mathbf{p})=\frac{\log\mathbf{p}}{\left\lceil\frac{K}{100}\cdot N\right\rceil}.$. Refer to Appendix˜A for more details on how other baselines are implemented.
LOS-Net can approximate GSFs and implement known baselines.  As the following result shows, our LOS-Net can, in fact, approximate virtually all GSFs of interest; intuitively, there exist sets of parameters such that it evaluates “arbitrarily close” to the target GSFs.
Proposition 2 (LOS-Net can approximate Equation˜4).
@@
To prove this result, we build on existing universality results on approximating continuous functions with Transformers [57], showing that our (generally non-continuous) target functions can be approximated by continuous functions. Importantly, Proposition˜2 implies that, as long as the LOS space of interest lies within a compact domain [^4] This is inherently satisfied when using probabilities; or via clamping in the case of logits or log-probs., our model can approximate the general GSF in Equation˜4 of LOSs for any LLM under mild conditions on $\kappa$, $T$, and $g$. Note that Proposition˜2 cannot be generally extended to $L_{\infty}$ due to the discontinuity of GSFs. The practical relevance of Proposition˜2, is underscored by the following: {corollary}[Approximation of Baselines by LOS-Net ] Our architecture, as defined in  Equation˜2, can arbitrarily well approximate, in the $L_{1}$ sense, any of the baseline methods in $\mathcal{B}$ when operating on context and token-vocabulary of, resp., maximal sizes $N_{\text{max}}$ and $V_{\text{max}}$. The above states that well-established, successful baselines (see class $\mathcal{B}$ in  Proposition˜1) can be approximated by LOS-Net. All proofs are enclosed in Appendix˜A.
@@
Act. Probe (incomp.) 73.00 $\pm$ 0.60 92.00 $\pm$ 1.00 72.00 $\pm$ 0.50 77.00 $\pm$ 0.50 81.00 $\pm$ 1.40 78.00 $\pm$ 0.40
Table 1: Test AUCs for HD over Mis-7b and L-3-8b (bold: best method, underlined: second best). orange indicates baselines requiring additional prompting/generations. Activation Probes, included as reference, are incomparable as they access model internals.
Figure 2: Transfer Test AUC to varying datasets (top, Mis-7b) and LLMs (bottom, IMDB fixed).
5 Experiments
We assess various aspects of learning with LOS via the following questions: (1) Is learning on LOS an effective approach for DCD and HD? Does it outperform baselines? And how important is $\mathbf{X}$, i.e., the TDS, in the pair $(\mathbf{X},\mathbf{p})$, often overlooked? (Sections˜5.2 and 5.1); (2) Does our model exhibit transfer capabilities across LLMs and datasets? (Section˜5.3); (3) What is the practical runtime of our approach and how robust is it w.r.t. $K$ (Sections˜5.4 and 5.5)?
In the following, we present our main results, and refer to Appendix˜B for additional results and details.
General setup.  Our experiments focus on the two tasks of DCD and HD; in all results presented next, hyperparameter $K$ is always set to $1,000$, while its impact is discussed in Section˜5.5. Aligning with prior work, we use datasets and LLMs from [43, 58] for DCD and [37] for HD, where we also experiment with an additional LLM (Qwen-2.5-7b-Instruct [53]). Further details are in subsequent sections. As performance metric, we use the area under the ROC curve (AUC), a standard metric in this domain [37, 43, 58]. We run each experiment with three different random seeds (when applicable) and report the mean along with the standard deviation of the results. All LOS-Net experiments were conducted using PyTorch [38] on a single NVIDIA L-40 GPU.
Newly introduced learning-based baselines.  In addition to task-specific baselines, we also introduce two novel learning-based baselines to appreciate the contribution of the TDS: ATP+R-MLP, ATP+R-Transf.. Specifically, we ablate information about the TDS and only process the ATP and rank information with, resp., an MLP or Transformer backbone. Formal definitions are in Section˜B.4.
5.1 Hallucination Detection
Datasets and LLMs.  We adopt datasets from Orgad et al. [37], following the same setup: the objective is to predict whether an LLM-generated response to a given input prompt is correct or not. We choose three datasets spanning various domains and tasks: HotpotQA without context [54], IMDB sentiment analysis [32], roles in Movies [37]. Details regarding the annotation process, splits and dataset sizes are in Section˜B.5.1. As the target LLMs, coherently with Orgad et al. [37], we use Mistral-7b-instruct-v0.2 [23] (Mis-7b) and LlaMa3-8b-instruct [47] (L-3-8b), and further experiment with Qwen-2.5-7b-Instruct [53] (Q-2.5-7b).
HD Baselines.
  1. 1.
    Aggregated probabilities/logits [16, 24, 49, 20]. They simply operate mean/max/min pooling over the ATP to score LLM confidence for error detection. We refer to them as Logit/Probas-mean/min/max.
  2. 2.
    P(True)Kadavath et al. [24] found that LLMs show reasonable calibration in assessing their own correctness via additional querying.
  3. 3.
    Semantic Entropy [13, 25]: a popular technique resorting to additional generations and an auxiliary entailment model to assess the uncertainty of LLM’s responses at a semantic level, argued, in turn, to be predictive for correctness. Note that both this method and the above P(True) require additional prompting and/or generations, making their detection latency orders of magnitudes higher than other methods in comparison, refer to Section˜5.4.
  4. 4.
    Activation Probes [37, 3, 4] are linear classifiers fitted over the LLM’s internal activations. They operate in the more restrictive white-box setup, thus not directly comparable to LOS-Net. Still, they constitute relevant performance references. We probe the last generated token at the layer maximizing validation performance.
@@ 5.1 Hallucination Detection
Results.Table˜1 presents a comprehensive summary of results on Mis-7b and L-3-8b. These clearly demonstrate that LOS-Net outperforms all gray-box baselines across all six dataset/LLM combinations, often by a significant margin. These also include P(True) and Semantic Entropy, which use auxiliary prompts or generations. We highlight how, on the IMDB dataset, LOS-Net achieves an AUC improvement of around 31 units over the best of these baselines for Mis-7b and 17 over the best baseline for L-3-8b. Intriguingly, we note how LOS-Net outperforms even white-box Activation Probes in 2 out of 6 combinations, while performing similarly in 3 of them. Our results also indicate that ATP learning-based baselines consistently underperform compared to LOS-Net, underscoring the critical role of the TDS, $\mathbf{X}$. Our ATP-based learnable baselines still outperform non-learnable probability-based methods in most cases, suggesting that a learning approach relying exclusively on ATP can still be a viable solution in certain scenarios. Results on Q-2.5-7b are consistent with the above findings, and are deferred to Section˜B.8.
@@ 5.2 Data Contamination Detection
DCD is often framed as a Membership Inference Attack (MIA) [44, 33, 43]. A DC dataset $D=\{q_{i},y_{i}\}_{i=1}^{\ell}$ contains $\ell$} text samples, where $q_{i}$ represents the text and $y_{i}$, the target, indicates whether it was part of the training data or not.
Datasets and LLMs.  We use three datasets to assess DCD, specifically: WikiMIA-32 and WikiMIA-64 [43] (excerpts from Wikipedia articles), as well as BookMIA [43] (excerpts from books). Henceforth, due to space limitations, we will only discuss details and results related to the latter, while referring readers to Section˜B.9 for the former. In BookMIA, positive members correspond to books known to be well memorized by certain OpenAI models [12], or otherwise known to (partly) be in the pretraining corpus of other open-source LLMs [1]. Non-members include excerpts from books released after 2023, necessarily absent from the pretraining corpus of the last ones. This dataset allows us to test LOS-Net in a realistic scenario akin to copyright-infringement detection. In particular, contrary to previous works, we propose a novel split that ensures all excerpts from the same book always appear either in the training or test split (and never in both). Details are enclosed in Section˜B.5.2. We attack LLMs considered in [1]: LlaMa-13b/30b [47] (L-13b/30b), Pythia-6.9b/12b[5] (P-6.9b/12b).
@@
Table 2: Test AUCs on BookMIA. ‘P’: Pythia, ‘L’: LlaMa-1 (bold: best, underlined: second best, pink: reference-based).
Figure 3: BookMIA zero-shot AUC – bold, : outperforms, resp., ref-free and ref-based.
DCD Baselines.  The Loss approach [55] directly uses the loss value as the detection score. The Reference (Ref) method [10] calibrates the target LLM’s perplexity leveraging a similar reference model known or supposed not to have memorized text of interest [^5] For example for Pythia-12b, a valid reference LLM would be the smaller Pythia-70M.. Both Zlib and Lowercase [10] are also reference-based methods: they utilize zlib compression entropy and lowercased text perplexity as reference for normalization. Lastly, Min-K% [43] and Min-K%++ [58] are reference-free methods, which examine token probabilities and average a subset of the minimum token scores, or a function thereof, over the input. For these baselines, we select their hyperparameters by maximizing performance on the validation set(s).
Results.  Refer to Table˜2 for results on BookMIA. LOS-Net attains exceptional results, largely surpassing other reference-free approaches. Among these last ones, ours is the only method that can match or outperform even the reference-LLM-based baselines. Importantly, (even partial) access to the TDS reveals crucial to obtain such strong reference-free performance: our ATP-based learnable methods – which only process features for the actual sequence tokens – incur indeed significant performance drops. As for WikiMIA, while full results are enclosed in Section˜B.9 (Table˜6), we point out how LOS-Net consistently surpasses all baselines across all combinations of LLMs and datasets. We also report the second-best method is MinK%++, followed by MinK%, consistent with the findings in [58].
@@ 5.3 Generalization to Other LLMs and Datasets
Zero-shot Cross-LLM Generalization in DCD.  We assess our model’s ability to detect DC in target LLMs that were unseen during training. Using the BookMIA benchmark and the setup described in Section˜5.2, we evaluate our model directly across different LLMs without any fine-tuning. This setup is particularly relevant in DCD scenarios where contamination information is not yet available for newly released LLMs. The results are presented in the heatmap shown in Figure˜3. We observe strong transferability: in 10/12 cases, our model achieves the best performance among reference-free approaches, highlighted in bold in Figure˜3. Interestingly, in 3/12 cases, LOS-Net (which is reference-free) even surpasses reference-based baselines, as indicated via a superscript of . We also observe particularly strong transfer across differently sized LLM architectures within the same family and highlight the surprising positive transfer from the largest LlaMa to Pythia models.
@@
Transfer Learning across LLMs and Datasets for HD.  Although LOS-Net delivered non-trivial generalization, its zero-shot application on HD was not sufficient to surpass the simpler probability-based techniques. This led us to investigate LOS-Net capabilities in a transfer learning setting. Specifically, we fix an LLM/dataset combination and fine-tune the corresponding pretrained LOS-Net either on the remaining LLMs for the same dataset, or the remaining datasets for the same LLM. All Test AUCs of our fine-tuned LOS-Net’s are in Figures˜4 and 5, Section˜B.6, while we report here two representative plots (see Figure˜2). On these heatmaps, superscript ‘’ indicates the fine-tuned LOS-Net is better than a counterpart trained from scratch in the same setting – testing for successful transfer; bold indicates it outperforms the best non-learnable probability-based method.
@@
Discussion.  First, LOS-Net exhibits solid transferability in both scenarios. The finetuned models consistently outperform their counterparts trained from scratch: 16/18 cases in both the cross-LLM (Figure˜4, Section˜B.6) and cross-dataset setups (Figure˜5, Section˜B.6) – see ‘’ on the off-diagonal entries. This highlights a generally positive transfer of LOS-Net’s learned representations across datasets and LLMs, and underscores the suitability of LOS as a data type in capturing generalizable patterns for HD. Second, from a practical perspective, we find that LOS-Net outperforms the best probability-based baseline in 15/18 cases for both the cross-LLM (Figure˜4) and cross-dataset (Figure˜5) scenarios – see bold on the off-diagonal entries. Focusing on the IMDB dataset, when training on L-3-8b and testing on Mis-7b (Figure˜2 (bottom)), our model substantially gains around 27 AUC units over the best probability-based baseline. This result underscores the possibility of transferring across LLMs. A similar trend is observed in the cross-dataset setup (Figure˜2 (top)): on Mis-7b, when training on HotpotQA or Movies and testing on IMDB, our model achieves a notable improvement of around 30 AUC units compared to the best baseline).